OpenStudy (anonymous):

Suppose y = 2x + 1 where x and y are functions of t. (a) If dx/dt = 6, find dy/dt when x = 4. (b) If dy/dt = 4, find dx/dt when x = 40.

OpenStudy (anonymous):

Can you do implicit differentiation?

OpenStudy (aum):

\[ y = 2x + 1 \\ \frac{dy}{dt} = 2\frac{dx}{dt} + 0 \\ a) \frac{dx}{dt} = 6. ~~\text{Sub above and find }\frac{dy}{dt} \]

OpenStudy (anonymous):

so then its 12?

OpenStudy (anonymous):

my online hw is saying thats the wrong answer

OpenStudy (adamaero):

with implicit diff. ...which is sad that I have to review at my level in Calculus

OpenStudy (aum):

\[ y = 2x + 1 \\ \frac{dy}{dx} = 2 = \large \frac{\frac{dy}{dt}}{\frac{dx}{dt}}\\ \frac{dy}{dt} = 2\frac{dx}{dt} \\ a) \frac{dx}{dt} = 6 \\ \frac{dy}{dt} = 2 * 6 = 12 \\ b) \frac{dy}{dt} = 4 \\ \frac{dx}{dt} = 4/2 = 2 \\ \]IDK why it is saying it is wrong.

OpenStudy (aum):

By any chance is \(\large y = 2^x + 1\) instead of \(\large y = 2x +1 \) ?